activity 2 solve the following problems. show all your works. 1 in a study of the life expectancy of 500 people in a certain geographic region, the mean age at death was 720 years, and the standard deviation was 5.3 years. if a sample of 50 people from this region is selected, find the probability that the mean life expectancy will be less than 70 years.

1
Expert's answer
2022-05-30T16:11:06-0400

Given "\mu=72.0" years, "\sigma=5.3" years, "N=500, n=50."


"\sigma_{X}=\dfrac{\sigma}{\sqrt{n}}\sqrt{\dfrac{N-n}{N-1}}=\dfrac{5.3}{\sqrt{50}}\sqrt{\dfrac{500-50}{500-1}}\approx0.7118"

"P(X<70)=P(Z<\dfrac{70-72}{0.7118})"

"\approx P(Z<-2.80985)\approx0.0025"